Two rugby players run towards each other and have an inelastic collision. Given the mass 70 kg and 60 kg respectively, and velocity 3 ms-1 and 4 ms-1. Calculate the final velocity and its direction of the respective players. Discuss how a rugby player’s bone develops if continuously performing the weight bearing exercise using Wolff’s Law and bone remodeling theory.

1
Expert's answer
2022-02-08T10:12:23-0500

Given:

"m_1=70\:\rm kg"

"m_2=60\:\rm kg"

"v_1=3\:\rm m/s"

"v_2=4\:\rm m/s"


The law of conservation of momentum says

"m_1v_1+m_2v_2=(m_1+m_2)u"

The speed of players

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}=\frac{70*3+60*4}{70+60}=3.5\:\rm m/s"

The velocity is directed toward direction of motion 60 kg player.